Save
Upgrade to remove ads
Busy. Please wait.
Log in with Clever
or

show password
Forgot Password?

Don't have an account?  Sign up 
Sign up using Clever
or

Username is available taken
show password


Make sure to remember your password. If you forget it there is no way for StudyStack to send you a reset link. You would need to create a new account.
Your email address is only used to allow you to reset your password. See our Privacy Policy and Terms of Service.


Already a StudyStack user? Log In

Reset Password
Enter the associated with your account, and we'll email you a link to reset your password.
focusNode
Didn't know it?
click below
 
Knew it?
click below
Don't Know
Remaining cards (0)
Know
0:00
Embed Code - If you would like this activity on your web page, copy the script below and paste it into your web page.

  Normal Size     Small Size show me how

Chapter 3

General Genetics- Exam 1

QuestionAnswer
Theories that have been proposed to explain the transmission of hereditary traits Theory of pangenesis Theory of preformationism
What is the blending theory? Factors that control hereditary traits are maleable They can blend together generation after generation
What experiment refuted blending theory? Gregor Mendel's pioneering experiments with garden peas refuted all of the above
Why was Mendel was successful when others were not? Mendel was successful when his predecessors were not because he chose a suitable organism, took time to ensure pure breeding, concentrated on one or a few discontinuous characters at a time, adopted quantitative form of analysis
Describe Mendel's Experimental System Used pure-breeding lines of garden peas Selected 7 traits to study, each of which had two distinguishable appearances He carefully designed his experiments and gathered mathematical data
What is cross fertilization? Pollen and egg are derived from different plants
What is self-fertilization? Pollen and egg are derived from the same plant
What is a monohybrid cross? A single-factor cross
What is the parental generation in Mendel's experiment? Pure bred lines
What is the first filial generation in Mendel's experiment? Offspring of cross-fertilization of parentals
What is the second filial generation in Mendel's experiment? Offspring of self-fertilization of F1 plants
What is the third filial generation in Mendel's experiment? Offspring of self-fertilization of F2 plants
What was the result of Mendel's experiment? 1. The F1 generation showed only one of the two parental traits 2. The F2 generation showed an ~ 3:1 ratio of the two parental traits 3. Reciprocal crosses yielded similar results
What did Mendel's results refute? Refuted a blending mechanism of heredity The data suggested a particulate theory of inheritance
What did Mendel postulate? A pea plant contains two discrete hereditary factors, one from each parent these factors (alleles) The two factors may be identical or different Segregation Random fertilization
Define homozygous When the alleles are identical
Define heterozygous When the alleles are different
What happens when the two factors differ in the Mendel's results? The dominant are expressed while recessive hidden
What is segregation? During gamete formation, the factors separate randomly, so half the gametes receive each factor
What is random fertilization? Male and female gametes randomly unite, reestablishing the paired factors
What is a genotype? The specific allelic composition of an individual With respect to plant height, a pea plant may have 3 different genotypes
What is a phenotype? The physical appearance of an individual With respect to plant height, a pea plant may have 2 different phenotypes
What is an example of different phenotypes that can be present? Tall; dwarf
The phenotype is a product of what? The phenotype is a product of the genotype and the environment
What is a punnett square A grid that enables one to calculate the results of simple genetics crosses
What is the genotypic ratio? 1: 2: 1
What is the phenotypic ratio? 3: 1
How did Mendel verify his principles of segregation? By performing a test-cross
What happens in a test-cross? In a test-cross an organism with known phenotype but unknown genotype is crossed with an organism that has the recessive phenotype (and hence is homozygous recessive)
What is a dihybrid cross? A two-factor cross
What is the law of independent assortment? During gamete formation, the segregation of any pair of hereditary determinants is independent of the segregation of other pairs
How was the law of independent assortment verified? By performing a test cross
What are modern geneticists often interested in? Modern geneticists are often interested in the relationship between the molecular expression of genes and the outcome of traits
What approach do modern geneticists use? Identify an individual with a defective copy of the gene Observe how this copy will affect the phenotype of the organism
What are loss-of-function alleles? The defective copies of genes Commonly inherited in a recessive manner
What is alternative method to the Punnet square? The forked line method
Steps to the fork line generation Treat each gene separately Consider phenotypic (or genotypic) differences
If n= # of heterozygous gene pairs involved Number of possible parental gametes= 2^n Number of different genotypes in zygotes= 3^n Number of different phenotypes in zygotes= 2^n Proportion of homozygous recessive in F2= 1/4^n
Propability formula/range # of favorable cases/total number of cases Ranges from 0 - 1
If an even has a probability of p, what is the probability of all alternative events q
Rules for combing probabilite Product rule Sum rule Binomial expansion equation (binomial theorem)
What is the product rule? The probability of the occurrence of independent events is the product of their respective probabilities
Probability of rolling a 3 AND then a 4 with a die? P(3 and 4) = P(3) X P(4)= 1/6 X 1/6= 1/36
What is the sum rule? The probability of the occurrence of one of several mutually exclusive events is the sum of their respective probabilities
Probability of rolling a 3 OR a 4 with a die? P(3 or 4) = P(3) + P(4) = 1/6 + 1/6 = 1/3
If the 2 events are not mutually exclusive, then.... (related to sum rule) P(A or B) = P(A) + P(B) - [P(A) X P(B)]
Probability of drawing an ace or a heart from a deck of cards P(A or ♥) = P(A) + P(♥) – [P(A) X P(♥)] = 4/52 + 13/52 – (4/52 X 13/52) = 17/52 – (1/52) = 16/52
What is the binomial expansion equation (definition)? The probability of the occurrence of unordered mutually exclusive events is defined by the binomial theorem
What is the binomial expansion equation (formula)? P = ((n!)/(x!(n-x)))p^x(q^n-x) where p = probability that the unordered number of events will occur n = total number of events x = number of events in one category p = individual probability of x q = individual probability of the other category
Two heterozygous brown-eyed (Bb) individuals have five children. What is the probability that two of the couple’s five children will have blue eyes? 26% of the time, a heterozygous couple’s five children will contain two with blue eyes and three with brown eyes
Why do we use chi-square analysis? Independent assortment/random fertilization are chance events, they are subject to random fluctuations To minimize the possibility that chance deviation will significantly alter the outcome
What is the chi-square analysis test used for? The chi-square test is a statistical method used to determine “goodness of fit” This refers to how close the observed data are to those predicted from a hypothesis
Created by: KristenR2025
 

 



Voices

Use these flashcards to help memorize information. Look at the large card and try to recall what is on the other side. Then click the card to flip it. If you knew the answer, click the green Know box. Otherwise, click the red Don't know box.

When you've placed seven or more cards in the Don't know box, click "retry" to try those cards again.

If you've accidentally put the card in the wrong box, just click on the card to take it out of the box.

You can also use your keyboard to move the cards as follows:

If you are logged in to your account, this website will remember which cards you know and don't know so that they are in the same box the next time you log in.

When you need a break, try one of the other activities listed below the flashcards like Matching, Snowman, or Hungry Bug. Although it may feel like you're playing a game, your brain is still making more connections with the information to help you out.

To see how well you know the information, try the Quiz or Test activity.

Pass complete!
"Know" box contains:
Time elapsed:
Retries:
restart all cards